Class 6 Maths Question 6.
Write the predecessor of:
(a) 94
(b) 10000
(c) 208090
(d) 7654321

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) Predecessor of 94 is 94 – 1 = 93 Hence, predecessor of 94 is 93. (b) Predecessor of 1000 is 10000 – 1 = 9999 Hence, predecessor of 10000 is 9999. (c) Predecessor of 208090 is 208090 -1 = 208089 Hence, predecessor of 208090 is 208089. (d) Predecessor of 7654321 is 7654321 – 1 = 7654320 Hence, predecessor of 7654321 is 7654320.